Get Help

Online: Request Help
Phone: (707) 826-HELP (4357)
Email: help@humboldt [dot] edu
In Person: Library 120 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
Campus Data Network Jul 8 2013 - 14:00 Network Users in Hadley House only *** Systems were rebooted and Service has been restored. *** TNS is investigating. Check back for updates.
Other Jul 5 2013 - 16:45 Wireless and ResNet Users The ResNet/Wireless DHCP server will be migrated from running on obsolete hardware to a Virtual Machine. The DHCP Server has been migrated.
Campus Phone System Jul 3 2013 - 15:20 Some campus voicemail users TNS is still investigating. We will know more after updates run Thursday morning TNS is investigating. Check back for updates. *** Update Service was restarted. Problem has cleared. A scheduled upgrade of the application will be upcoming to resolve the issue.
CMS/Peoplesoft Jul 3 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following items will be installed: MP 9.0, HR retrofit, and Self Service retrofit.
Campus Data Network Jun 28 2013 - 09:21 Campus Users Attempting to Connect to Sites Off-Campus ** UPDATE Sunday 6/30 The main network link has been restored as well **** (12:31) redundant link is back up. Network connectivity restored. Check back for updates on the main campus network link ** Still No ETA -- Check back for updates ** We're in touch with CENIC currently and they are investigating. Check back for updates. We are back !!!
Off-Campus Internet Connection Jun 30 2013 - 23:00 Any network user trying to connect off campus Due to the Network Outage on Friday 6/28, maintenance needs to be performed. There should not be any affect to the campus, but there could be rolling outages during this time perioud. If you are working at this time, please save your work frequently.
Other Jun 27 2013 - 17:00 Students, Faculty, Staff Migration to new server. System is Available.
Other Jun 27 2013 - 17:00 All Users The aging hardware serving the Campus DHCP server (Except Housing) will be retired.
Google Apps - Mail Jun 26 2013 - 11:30 All Google Mail Users Help Desk confirmed this issue to be campus-wide. Opening a ticket with Google. Google opened Case #03233237, and confirmed that a lot of their customers were having the same problem. The problem has allegedly been fixed in the meantime.
Moodle Jul 21 2013 - 05:00 Moodle Users We will be installing a critical patch.
Moodle Jun 28 2013 - 20:00 ** THIS EVENT HAS BEEN CANCELLED ** Moodle Users ** THIS EVENT HAS BEEN CANCELLED -- WE WILL BE RESCHEDULING THIS MAINTENANCE ** Moodle will be down for patching from Friday 8:00pm to 8:00am. This patching is required to fix a critical bug.
Campus Data Network Jun 24 2013 - 13:15 Network Users at Natural History Museum only *** UPDATE Power has been restored in town and so the network service is back up and working *** -- TNS investigated and found there is a power outage in town. Unknown at this time how long the power will be out. -- Check back for updates.
Campus Data Network Jun 21 2013 - 12:53 Marine Lab network users only *** UPDATE - Service has been restored, TNS is still investigating the cause *** TNS is investigating. Check back for updates.
Campus Data Network Jun 19 2013 - 16:38 All Hadley House network users *** UPDATE - Service has been restored. *** TNS is investigating. Check back for updates.
CMS/Peoplesoft Jun 19 2013 - 18:00 All CFS Production (FCFSPRD) and CFS Data Warehouse application and database services users. During our regularly scheduled maintenance window the following items will be installed: PeopleSoft Fixes for 9.0 AP, AR Aging by Customer Report, Enterprise Fees, Instructional Equipment, Reporting Solution, and GAAP Reports.
Other Jun 18 2013 - 09:41 All Portal users HSU Portal server was temporarily not accessible for approximately 10 minutes. It is now back up and available.
Other Jun 17 2013 - 15:13 Sophos Home Users Sophos Home Users may have noticed in the past few days that there was a red "x" on the Sophos icon in their system trays indicating that Sophos could not contact the server. The system has been restarted and updating service is restored.
Other Jun 15 2013 - 13:00 All Nolij Users / OAK database server users Shutdown delays due to excessive Database server Memory Use.
Other Jun 15 2013 - 05:00 All Users Scheduled Maintenance
Off-Campus Internet Connection Jun 6 2013 - 09:03 All campus community ATT issue identified affecting multiple sites in Northern California. CENIC is working with ATT to resolve the issue.
Moodle Jun 8 2013 - 12:00 >> This Event Has Been Cancelled << Faculty & Students >> This Event Has Been Cancelled << Maintenance window.
CMS/Peoplesoft Jun 6 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During an additional planned outage outside of our regularly scheduled maintenance window, the following items will be installed: DARS Interactive Changes and CAS Authentication. Special Notes related to CAS Authentication: 1. You may see the login page flash on your browser, before you are authenticated. 2. If you access HCM production without going through the Portal, you will now be redirected to the CAS login. 3. If you have a batch account, you will receive a separate login URL after the changes have been successfuly verified.
Domain Name Service Jun 5 2013 - 17:00 All Users Rescanning the RAID controller after adding a hot-spare drive unexpectedly led to a system state requiring a reboot. The server has been rebooted, everything looks well.
Other Jun 5 2013 - 14:30 Virtual Desktop Users (not UPD) Network reconfiguration
Other Jun 5 2013 - 09:24 Users of OBI and SDR system
CMS/Peoplesoft Jun 5 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly schedule maintenance window the following item will be installed: Oracle 11g Upgrade. NOTE: The Absence Hint Optimization baseline change will go in on 6/6 at 10:00 a.m with no system outage.
Other May 30 2013 - 15:59 All campus users ITS is investigating, stay tuned. [Connectivity issues resolved-still investigating cause.]
CMS/Peoplesoft May 29 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following items will be installed: MP 8.05, MP 8.06, MP 8.07, MP 8.08, MP 8.09, Finance Aid Self-Service Retrofit, Orientation Modification View Change, and Absence Search Improvement.
Campus Phone System May 21 2013 - 10:55 All campus phone users making off-campus local phone calls *** UPDATE - A power surge at the AT&T Central Office caused the problem, the circuit has been replaced and service is restored *** AT&T is looking into it. Please check back for updates.
Campus Phone System May 20 2013 - 08:10 All campus voicemail users *** UPDATE - System is working now, TNS is still investigating cause and monitoring system *** Unknown at this time. TNS is investigating. Check back for updates.
feedback